what is the equation for a line with these points (0, -1), (2, -5), (4, -9)

Answer by lwsshak3(11628)

what is the equation for a line with these points (0, -1), (2, -5), (4, -9)
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
From point (0,-1), y-intercept,b,=-1
slope=∆y/∆x=(-5-(-1))/2-0)=-4/2=-2
equation of straight line: y=-2x-1